# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

### 05 Mar 2026 v1

@@ II.5 Maximum attainable energy during cosmological magnetogenesis
Fig. 3 shows the redshift evolution of the maximum attainable proton energy for representative values of the turbulence parameter $\epsilon$. Evaluated near the turn-on epoch $z_{\rm on}\sim 1.7$, where $B(z_{\rm on})\sim 10^{-9}\,\mathrm{G}$ in our fiducial model, Eq. (28) implies that proton energies can, under optimistic optimistic assumptions corresponding to the strong-scattering limit, reach values up to $\sim 10^{2}\,\mathrm{GeV}$. The strong dependence $E_{\max}\propto\epsilon^{2}$ implies that larger turbulence amplitudes lead to higher maximum energies, while the overall redshift dependence remains unchanged. At higher redshift ($z\gtrsim z_{\rm on}$), the magnetic field remains weak and the Hubble time short, so that $E_{\max}$ is strongly suppressed and remains far below the relativistic regime. Thus, even under optimistic assumptions, cosmological stochastic acceleration cannot generate a substantial high-energy particle population prior to the onset of nonlinear structure formation. We emphasize that Eq. (28) provides an upper bound. Additional physical effects—including Coulomb and ionization losses, incomplete turbulence development, or deviations from Bohm-like diffusion—would further reduce the achievable energy. Consequently, while stochastic acceleration may in principle produce $\mathcal{O}(10$-$10^{2})\,\mathrm{GeV}$ protons near $z_{\rm on}\sim 1.7$, the emergence of structure-formation shocks remains a far more efficient and robust mechanism for accelerating CRs to higher energies.
III Pre-Structure Stochastic Acceleration in the Early Universe
@@
To quantify the magnitude of any seed nonthermal population produced by the instability-assisted stochastic process discussed above, we introduce a transport model for the isotropic proton distribution function. The goal is not to construct a fully self-consistent cosmological CR population synthesis, but rather to translate the acceleration-time constraints into an explicit estimate of the resulting momentum-space spectrum and energy content.
III.1 Isotropic Fokker-Planck equation in an expanding background
@@
Let $f(p,z)$ be the isotropic phase-space distribution of protons, normalized such that the differential number density is $dn=4\pi p^{2}f(p,z)\,dp$. Neglecting spatial gradients and assuming isotropic pitch-angle scattering, the momentum-space Fokker-Planck equation reads
@@
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)-\frac{f}{t_{\rm esc}(p,z)},\,$ (30)
where $D_{pp}$ is the momentum diffusion coefficient associated with second-order Fermi acceleration, $\dot{p}_{\rm loss}$ represents systematic momentum losses, $Q$ is a source term (e.g., injection from the thermal pool), and $t_{\rm esc}$ is an effective escape time from the acceleration region.
Cosmological expansion is incorporated by transforming from $t$ to redshift $z$ using
$\frac{dt}{dz}=-\frac{1}{(1+z)H(z)}\,$, (31)
with $H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}$. Eq. (30) then becomes
@@
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)\,.$ (32)
@@
$D_{pp}(p,z)=\frac{p^{2}}{t_{\rm acc}(z)}~(p\gtrsim p_{\rm inj}),$ (33)
corresponding to a momentum diffusion coefficient proportional to $p^{2}$, with a redshift-dependent normalization determined by the microphysical scattering rate. Here, $p_{\rm inj}$ denotes the injection momentum for stochastic acceleration. This form represents the minimal choice consistent with the timescale analysis presented in Section II.
In the present model, we assume that the escape timescale is much longer than the characteristic acceleration and loss timescales, $t_{\rm esc}\gg t_{\rm acc},~t_{\rm loss}$. Here, the characteristic loss timescale is defined as $t_{\rm loss}\equiv p/\dot{p}_{\rm loss}$. Cosmological expansion inevitably induces adiabatic momentum losses,
$\dot{p}_{\rm ad}(p,z)=-H(z)\,p\,.$ (34)
In principle, additional Coulomb losses can be included at sub-relativistic energies, $\dot{p}_{\rm loss}=\dot{p}_{\rm ad}+\dot{p}_{\rm C}$, to assess whether any incipient suprathermal tail is efficiently thermalized. However, since our main interest lies in the relativistic tail near and above the turn-on epoch, the adiabatic term already captures the unavoidable cosmological sink. Therefore, we adopt $\dot{p}_{\rm loss}\approx\dot{p}_{\rm ad}$. With Eqs. (33) and (34), Eq. (32) becomes
$-(1+z)H(z)\,\frac{\partial f}{\partial z}$ $\approx\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\frac{p^{2}}{t_{\rm acc}(z)}\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\left(-H(z)p\right)f\right]+Q(p,z)\,.$ (35)
Figure 4: Proton spectra $p^{4}f(p,z)$ from the Fokker-Planck model for $z=10,5,2$ and $1.7$. Panel (a): $\epsilon=0.05$; panel (b): $\epsilon=0.2$. The initial distribution at $z=10$ is Maxwellian with $k_{B}T_{0}=0.86~{\rm eV}$.
Figure 5: Same as Fig. 4, but including a weak pre-existing suprathermal power-law tail ($s=4.2$) in the initial distribution.
III.2 Numerical implementation and early-universe heating
To assess the magnitude of stochastic heating prior to the magnetic-field turn-on epoch, we solve Eq. (35) numerically in the redshift interval $z=10\rightarrow z_{\rm on}$, where the magnetic field remains well below its rapid-amplification regime. In this early phase, the goal is to determine whether the instability-assisted second-order Fermi process can produce a significant seed nonthermal component over a Hubble time.
@@ III.2 Numerical implementation and early-universe heating
$f_{\rm MW}(p,z=10)=\frac{n}{(2\pi m_{p}k_{B}T_{0})^{3/2}}\exp\!\left(-\frac{p^{2}}{2m_{p}k_{B}T_{0}}\right),$ (36)
where $n$ denotes the proton number density and $k_{B}T_{0}=0.86~{\rm eV}$. No explicit source term is included ($Q=0$), and escape is neglected, so that the evolution is governed solely by stochastic momentum diffusion and adiabatic losses. The magnetic field $B(z)$ is mapped to redshift using the cosmological relation $dt/dz=-[(1+z)H(z)]^{-1}$. In the early regime ($z\gtrsim z_{\rm on}$), $B(z)$ evolves smoothly and remains far below the explosive growth associated with $z<z_{\rm on}$, ensuring that $t_{\rm acc}(z)$ varies gradually.
Eq. (35) is solved on a logarithmic momentum grid using a fully implicit scheme in redshift. The equation is written in conservative flux form in momentum space,
$\mathcal{R}(f)=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}\frac{\partial f}{\partial p}-p^{2}\dot{p}_{\rm ad}f\right],$ (37)
@@
Fig. 4 shows the resulting spectra $p^{4}f(p,z)$ at $z=10,5,$ and $2$ for two representative turbulence levels, $\epsilon=0.05$ and $\epsilon=0.2$. In both cases, the thermal peak shifts toward lower momentum as redshift decreases. This behavior reflects adiabatic cooling, for which $p\propto a^{-1}\propto(1+z)$. The corresponding reduction in peak amplitude is likewise consistent with the decreasing thermal energy density of the expanding background. The stochastic term introduces a modest broadening of the high-momentum tail relative to pure adiabatic evolution. For $\epsilon=0.05$, the acceleration time remains long compared to the Hubble time throughout $z=10\rightarrow 2$, and the distribution remains close to Maxwellian. For $\epsilon=0.2$, the diffusion coefficient increases by a factor of four, leading to a slightly enhanced suprathermal component; however, adiabatic cooling still dominates the overall evolution over this redshift interval.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the situation begins to change. The amplification of the magnetic field reduces the stochastic acceleration time $t_{\rm acc}$, allowing the diffusion term to partially compete with the adiabatic loss term. As a result, a mild suprathermal tail begins to emerge in the particle spectrum. Although the nonthermal component remains energetically subdominant, this behavior indicates that stochastic acceleration becomes progressively more effective near the onset of magnetic-field amplification.
We do not extend the mini Fokker–Planck integration below $z_{\rm on}$, since in this regime large-scale structure formation shocks are expected to dominate particle acceleration. The present model is therefore designed to quantify the pre-shock stochastic heating phase and the magnitude of any seed suprathermal population prior to the onset of structure formation. In the next section, we compare the characteristic acceleration timescale of the instability-assisted stochastic process with that of DSA in the Bohm-diffusion limit, in order to determine which mechanism governs CR production once shocks emerge.
III.3 Sensitivity to a pre-existing suprathermal population
The calculations above assume that the proton distribution at $z=10$ is purely Maxwellian. However, weak nonthermal particle populations may already exist in the early universe as a result of earlier energetic events, such as Population III supernova explosions or structure-formation shocks [e.g., 23].

### 01 Apr 2026 v2

@@ 2.5 Maximum attainable energy during cosmological magnetogenesis
Figure 3 shows the redshift evolution of the maximum attainable ion energy for representative values of the turbulence parameter $\epsilon$. Evaluated near the turn-on epoch $z_{\rm on}\sim 1.7$, where $B(z_{\rm on})\sim 10^{-9}\,\mathrm{G}$ in our fiducial model, Eq. (2.31) implies that ion energies can, under optimistic optimistic assumptions corresponding to the strong-scattering limit, reach values up to $\sim 10^{2}\,\mathrm{GeV}$. The strong dependence $E_{\max}\propto\epsilon^{2}$ implies that larger turbulence amplitudes lead to higher maximum energies, while the overall redshift dependence remains unchanged. At higher redshift ($z\gtrsim z_{\rm on}$), the magnetic field remains weak and the Hubble time short, so that $E_{\max}$ is strongly suppressed and remains far below the relativistic regime. Thus, even under optimistic assumptions, cosmological stochastic acceleration cannot generate a substantial high-energy particle population prior to the onset of nonlinear structure formation. We emphasize that Eq. (2.31) provides an upper bound. Additional physical effects—including Coulomb and ionization losses, incomplete turbulence development, or deviations from Bohm-like diffusion—would further reduce the achievable energy. Consequently, while stochastic acceleration may in principle produce $\mathcal{O}(10$-$10^{2})\,\mathrm{GeV}$ ions near $z_{\rm on}\sim 1.7$, the emergence of structure-formation shocks remains a far more efficient and robust mechanism for accelerating CRs to higher energies.
3 Refined Constraints Including Coulomb Losses in a Fully Ionized Medium
In Section 2, we derived the conditions for stochastic acceleration by considering adiabatic losses associated with cosmological expansion. This provides an optimistic estimate of the acceleration efficiency, corresponding to a lower bound on the acceleration time and a minimal requirement for particle energization. In this section, we refine the analysis by including Coulomb losses, which arise from interactions with background electrons and become important at sub-relativistic energies. We emphasize that this effect is distinct from the role of Coulomb collisions in pitch-angle scattering: while Coulomb scattering is negligible compared to instability-driven scattering (i.e., $\nu_{ii}\ll\nu_{\rm scatt}$), Coulomb interactions can still dominate the energy evolution of particles through collisional energy losses, in the sense that the Coulomb loss rate can exceed the stochastic acceleration rate ($t_{C}^{-1}\gtrsim t_{\rm acc}^{-1}$). We focus on a fully ionized medium, appropriate for the post-reionization intergalactic medium. Since the turn-on redshift derived in this work lies at $z_{\rm on}\sim 1.7$, well below the epoch of reionization ($z\sim 6$), this assumption is well justified. In such an environment, Coulomb interactions dominate the collisional energy losses, allowing us to isolate the impact of collisional cooling on stochastic acceleration in a physically well-defined and tractable regime. We note that additional ionization losses in partially neutral regions would further shorten the cooling timescale and thus strengthen the constraints derived below.
3.1 Energy loss processes in the pre-structure intergalactic medium
In addition to adiabatic losses due to cosmological expansion, charged particles in the pre-structure intergalactic medium experience energy losses through Coulomb interactions with free electrons and, in partially neutral regions, through ionization processes. The characteristic energy loss timescales are defined as
$t_{\rm C}(E,z)\equiv\frac{E}{|\dot{E}_{\rm C}(E,z)|},\qquad t_{\rm ioniz}(E,z)\equiv\frac{E}{|\dot{E}_{\rm ioniz}(E,z)|}.$ (3.1)
Both processes are most effective at sub-relativistic energies and scale inversely with the ambient density, $t_{\rm C},t_{\rm ioniz}\propto n^{-1}(z)$. Since the baryon density evolves as $n(z)\propto(1+z)^{3}$, the corresponding cooling times decrease rapidly toward higher redshift. In ionized regions of the intergalactic medium, Coulomb interactions with free electrons dominate the collisional energy losses, while ionization losses become important only in partially neutral environments. In the following, we adopt a fully ionized medium as a fiducial model, appropriate for the post-reionization universe, and therefore focus on Coulomb losses as the primary non-adiabatic cooling process. Coulomb interactions of ions are generally dominated by scattering off thermal electrons [e.g., 35], owing to the cumulative effect of frequent small-angle encounters with light particles. Accordingly, we approximate the Coulomb energy loss rate using the electron contribution.
For sub-relativistic and mildly relativistic ions in an ionized plasma, the Coulomb energy loss rate is dominated by interactions with background electrons and can be approximated as
$\dot{E}_{\rm C}(E,z)\approx\frac{4\pi e^{4}n_{e}(z)\ln\Lambda}{m_{e}v(E)},$ (3.2)
where $n_{e}(z)$ is the electron number density, $\ln\Lambda$ is the Coulomb logarithm, and $v(E)$ is the particle velocity. This leads to a characteristic Coulomb loss timescale
$t_{\rm C}(E,z)\approx\frac{Em_{e}v(E)}{4\pi e^{4}n_{e}(z)\ln\Lambda}.$ (3.3)
This expression explicitly shows that Coulomb losses become increasingly efficient at low energies due to the inverse dependence on particle velocity.
The total effective loss rate is given by
$t_{\rm loss}^{-1}(E,z)=t_{H}^{-1}(z)+t_{C}^{-1}(E,z),$ (3.4)
so that in the limits $t_{C}\ll t_{H}$ or $t_{H}\ll t_{C}$, the effective loss timescale reduces to the shorter of the two. We note that the inclusion of ionization losses in partially neutral regions would further shorten the cooling timescale at low energies, thereby strengthening the constraints derived below.
3.2 Refined critical magnetic field
Including these additional loss processes, the condition for stochastic acceleration becomes
$t_{\rm acc}(E,z)<t_{\rm loss}(E,z).$ (3.5)
This leads to a refined critical magnetic field, $B_{\rm crit}^{\rm eff}(E,z)$, defined implicitly by
$t_{\rm acc}(E,z)=t_{\rm loss}(E,z).$ (3.6)
Using the optimistic acceleration timescale adopted in Section 2, we obtain an explicit expression for the refined critical field,
$B_{\rm crit}^{\rm eff}(E,z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{1}{\Delta_{\rm ex}^{3/2}t_{\rm loss}(E,z)}.$ (3.7)
This expression reduces to the Hubble-limited critical field when $t_{\rm loss}=t_{H}$,
$B_{\rm crit}(z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (3.8)
More generally, the two are related by
$B_{\rm crit}^{\rm eff}(E,z)=\frac{B_{\rm crit}(z)}{H(z)t_{\rm loss}(E,z)}.$ (3.9)
Figure 4: Redshift evolution of the magnetic field $B(z)$ (black) compared with the refined critical field $B_{\rm crit}^{\rm eff}(E,z)$ for different particle energies including Coulomb losses. The gray curve shows the Hubble-limited result. Coulomb losses strongly suppress acceleration at low energies, while at high energies $B_{\rm crit}^{\rm eff}$ converges to the Hubble-limited case, indicating an effective energy threshold for stochastic acceleration.
Since Coulomb losses reduce the effective loss timescale at low energies, $t_{\rm loss}(E,z)\leq t_{H}(z)$, the refined critical field satisfies
$B_{\rm crit}^{\rm eff}(E,z)\gtrsim B_{\rm crit}(z),$ (3.10)
indicating that a stronger magnetic field is required for efficient acceleration when realistic energy losses are taken into account. In particular, in the regime where Coulomb losses dominate ($t_{\rm loss}\approx t_{\rm C}$), the critical field becomes explicitly energy dependent,
$B_{\rm crit}^{\rm eff}(E,z)\propto\frac{1}{v_{\rm tur}^{2}(z)}\frac{v(E)}{t_{\rm C}(E,z)\Delta_{\rm ex}^{3/2}},$ (3.11)
reflecting the fact that low-energy particles are more strongly affected by collisional cooling. This introduces an additional suppression of stochastic acceleration compared to the adiabatic-only estimate.
Figure 4 shows the redshift evolution of the refined critical magnetic field $B_{\rm crit}^{\rm eff}(E,z)$ for a range of particle energies, together with the magnetic-field evolution $B(z)$ and the Hubble-limited critical field. At low energies, Coulomb losses significantly reduce the effective loss timescale ($t_{\rm loss}\approx t_{\rm C}$), leading to a substantial increase in $B_{\rm crit}^{\rm eff}$. As a result, the required magnetic field for efficient stochastic acceleration becomes much larger than in the adiabatic-only case, and the crossing with $B(z)$ is either shifted to much lower redshift or entirely suppressed. This demonstrates that collisional cooling strongly inhibits particle acceleration near the injection scale. In contrast, at sufficiently high energies where $t_{\rm C}\gg t_{H}$, the effective loss timescale is governed by cosmological expansion ($t_{\rm loss}\approx t_{H}$), and $B_{\rm crit}^{\rm eff}$ converges to the Hubble-limited critical field derived in Section II. In this regime, stochastic acceleration becomes largely insensitive to Coulomb losses, and the corresponding turn-on redshift approaches the relativistic baseline shown in Figure 1. These results clearly demonstrate that Coulomb losses introduce a strong energy dependence in the conditions for stochastic acceleration, primarily affecting low-energy particles while leaving high-energy particles largely unaffected.
The comparison between $B(z)$ and $B_{\rm crit}^{\rm eff}(E,z)$ further implies the existence of an effective low-energy threshold for stochastic acceleration. Below this threshold, Coulomb losses dominate over acceleration, preventing particles from gaining energy efficiently and suppressing their injection into the acceleration process. In the fiducial model considered here, this threshold is found to lie at approximately the ${\mathcal{O}}(10)$ keV level. Particles with energies below this scale experience rapid collisional cooling, such that $t_{\rm C}\ll t_{\rm acc}$ over the relevant redshift range. As a result, stochastic acceleration is effectively quenched. Conversely, for particles with energies above this threshold, the Coulomb loss timescale increases rapidly, allowing stochastic acceleration to operate and eventually approach the Hubble-limited regime. We therefore conclude that Coulomb losses impose an effective injection energy for stochastic acceleration, transforming the acceleration condition from a purely cosmological constraint into an energy-dependent threshold problem. At sufficiently high energies, however, where $t_{\rm C}\gg t_{H}$, the acceleration process becomes limited by cosmological expansion, and the maximum attainable energy remains essentially unchanged from the adiabatic-only case.
4 Pre-Structure Stochastic Acceleration in the Early Universe
@@
To quantify the magnitude of any seed nonthermal population produced by the instability-assisted stochastic process discussed above, we introduce a transport model for the isotropic ion distribution function. The goal is not to construct a fully self-consistent cosmological CR population synthesis, but rather to translate the acceleration-time constraints into an explicit estimate of the resulting momentum-space spectrum and energy content.
4.1 Isotropic Fokker-Planck equation in an expanding background
@@
Let $f(p,z)$ be the isotropic phase-space distribution of ions, normalized such that the differential number density is $dn=4\pi p^{2}f(p,z)\,dp$. Neglecting spatial gradients and assuming isotropic pitch-angle scattering, the momentum-space Fokker-Planck equation reads
@@
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)-\frac{f}{t_{\rm esc}(p,z)},\,$ (4.1)
where $D_{pp}$ is the momentum diffusion coefficient associated with second-order Fermi acceleration, $\dot{p}_{\rm loss}$ represents systematic momentum losses, $Q$ is a source term (e.g., injection from the thermal pool), and $t_{\rm esc}$ is an effective escape time from the acceleration region. Cosmological expansion is incorporated by transforming from $t$ to redshift $z$ using
$\frac{dt}{dz}=-\frac{1}{(1+z)H(z)}\,$, (4.2)
with $H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}$. Eq. (4.1) then becomes
@@
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)\,.$ (4.3)
@@
$D_{pp}(p,z)=\frac{p^{2}}{t_{\rm acc}(z)}~(p\gtrsim p_{\rm inj}),$ (4.4)
corresponding to a momentum diffusion coefficient proportional to $p^{2}$, with a redshift-dependent normalization determined by the microphysical scattering rate. Here, $p_{\rm inj}$ denotes the injection momentum for stochastic acceleration. This form represents the minimal choice consistent with the timescale analysis presented in Section 2. In the present model, we assume that the escape timescale is much longer than the characteristic acceleration and loss timescales, $t_{\rm esc}\gg t_{\rm acc},~t_{\rm loss}$. The characteristic loss timescale is defined as $t_{\rm loss}\equiv p/\dot{p}_{\rm loss}$. Cosmological expansion inevitably induces adiabatic momentum losses,
$\dot{p}_{\rm ad}(p,z)\approx-H(z)p.$ (4.5)
In addition, particles experience Coulomb losses due to interactions with background electrons, which are particularly important at sub-relativistic energies. The total momentum loss rate is therefore given by
$\dot{p}_{\rm loss}(p,z)=\dot{p}_{\rm ad}(p,z)+\dot{p}_{\rm C}(p,z),$ (4.6)
where the Coulomb momentum loss rate can be approximated as
$\dot{p}_{\rm C}(p,z)\approx-\frac{4\pi e^{4}n_{e}(z)\ln\Lambda}{m_{e}v^{2}(p)}.$ (4.7)
Here $n_{e}(z)$ is the electron number density, $\ln\Lambda$ is the Coulomb logarithm, and $v(p)$ is the particle velocity. This expression shows that Coulomb losses increase rapidly toward low velocities, leading to efficient thermalization of particles near the injection scale. Including both contributions, Eq. (4.3) becomes
$-(1+z)H(z),\frac{\partial f}{\partial z}$ $\approx\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\frac{p^{2}}{t_{\rm acc}(z)}\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\left(\dot{p}_{\rm ad}(p,z)+\dot{p}_{\rm C}(p,z)\right)f\right]$
$+Q(p,z).$ (4.8)
The Coulomb term introduces an additional sink of particle momentum at low energies, which can suppress the formation of a suprathermal tail by efficiently thermalizing particles near the injection scale.
We note that adiabatic momentum losses due to cosmological expansion are applicable to particles residing in the unbound intergalactic medium considered in this work. In contrast, particles confined within gravitationally bound systems, such as galaxies or galaxy clusters, do not experience cosmological adiabatic cooling. Since our analysis focuses on the pre-structure phase prior to nonlinear collapse, the assumption $\dot{p}_{\rm ad}\approx-H(z)p$ remains valid throughout the redshift interval of interest.
4.2 Numerical implementation and early-universe heating
To assess the magnitude of stochastic heating prior to the magnetic-field turn-on epoch, we solve Eq. (4.8) numerically over the redshift interval $z=10\rightarrow z_{\rm on}$, where the magnetic field remains well below its rapid-amplification regime. In this early phase, the goal is to determine whether the instability-assisted second-order Fermi process can produce a significant seed nonthermal component over a Hubble time.
@@ 4.2 Numerical implementation and early-universe heating
$f_{\rm MW}(p,z=10)=\frac{n_{i}}{(2\pi m_{i}k_{B}T_{0})^{3/2}}\exp\left(-\frac{p^{2}}{2m_{i}k_{B}T_{0}}\right),$ (4.9)
where $n_{i}$ denotes the ion number density. We emphasize that this setup is not intended to represent the volume-averaged thermal state of the universe at $z=10$, but rather a simplified diffuse ionized medium used as a fiducial environment for assessing the competition between stochastic acceleration and cooling processes. Accordingly, $T_{0}$ should be interpreted as an effective model parameter. In the fiducial case, we adopt $k_{B}T_{0}=0.86~{\rm eV}$, corresponding to a moderately heated plasma. No explicit source term is included ($Q=0$), and escape is neglected, so that the evolution is governed solely by stochastic momentum diffusion and energy losses. The magnetic field $B(z)$ is mapped to redshift using the cosmological relation $dt/dz=-[(1+z)H(z)]^{-1}$. In the early regime ($z\gtrsim z_{\rm on}$), $B(z)$ evolves smoothly and remains far below the explosive growth associated with $z<z_{\rm on}$, ensuring that $t_{\rm acc}(z)$ varies gradually.
Eq. (4.8) is solved on a logarithmic momentum grid using a fully implicit scheme in redshift. The equation is written in conservative flux form in momentum space,
$\mathcal{R}(f)=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}\frac{\partial f}{\partial p}-p^{2}(\dot{p}_{\rm ad}+\dot{p}_{\rm C})f\right],$ (4.10)
@@
Figure 5: Ion spectra $p^{4}f(p,z)$ from the Fokker–Planck model at $z=10,5,2,$ and $1.7$. Panel (a): full solution including Coulomb and adiabatic losses; panel (b): adiabatic-loss-only case. The initial distribution at $z=10$ is a Maxwellian with $k_{B}T_{0}=0.86~{\rm eV}$. The turbulent velocity is parameterized as $v_{\rm tur}=\epsilon c_{s}$ with $\epsilon=0.2$. Coulomb losses suppress the formation of a suprathermal tail by efficiently cooling low-energy particles.
Figure 5 shows the evolution of the particle spectra $p^{4}f(p,z)$ at $z=10,5,2,$ and $1.7$, comparing the full solution including Coulomb losses (left panel) with the adiabatic-loss-only case (right panel). In both cases, the thermal peak shifts toward lower momentum as redshift decreases, reflecting adiabatic cooling for which $p\propto a^{-1}\propto(1+z)$. The corresponding decrease in peak amplitude is likewise consistent with the dilution of the thermal energy density in an expanding background. In the absence of Coulomb losses (right panel), stochastic momentum diffusion produces a modest broadening of the high-momentum tail relative to pure adiabatic evolution. For stronger turbulence, the suprathermal component becomes slightly more pronounced as the system approaches the magnetic-field turn-on epoch, indicating that stochastic acceleration begins to compete with adiabatic cooling. In contrast, when Coulomb losses are included (left panel), the low-energy population is efficiently thermalized, and the formation of a suprathermal tail is strongly suppressed over the entire redshift range considered. Even for relatively strong turbulence, the distribution remains close to a Maxwellian, with only minimal deviation at high momenta. This behavior reflects the fact that Coulomb cooling dominates over stochastic acceleration near the injection scale, preventing particles from being efficiently accelerated out of the thermal pool. As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the acceleration timescale decreases due to magnetic-field amplification. However, the results indicate that Coulomb losses continue to limit the development of a significant nonthermal component in the pre-shock phase. We therefore conclude that stochastic acceleration alone is unable to generate a substantial suprathermal seed population prior to the onset of structure-formation shocks.
We do not extend the Fokker–Planck integration below $z_{\rm on}$, since in this regime large-scale structure formation shocks are expected to dominate particle acceleration. The present model is therefore designed to quantify the pre-shock stochastic heating phase and to assess the extent to which a seed nonthermal population can be established before shock-driven acceleration takes over.
4.3 Limited impact of a pre-existing suprathermal population
While the preceding analysis assumes a purely thermal initial distribution, a weak suprathermal particle population may arise from a variety of processes operating in weakly magnetized plasmas, such as plasma instabilities, turbulence, or residual nonthermal components inherited from earlier evolutionary stages. Even a small nonthermal tail could in principle act as a seed population for subsequent acceleration.